OpenStudy (anonymous):

Two tugboats that are a = 130 ft apart pull a barge, as shown. If the length of one cable is b = 218 ft and the length of the other is c = 230 ft, find the angle formed by the two cables.

hero (hero):



hero (hero):

Use law of cosines

OpenStudy (anonymous):

yea i know i have to use law of cosines but i forgot my book in school and google confused me. do you mind if u solve it for me please

OpenStudy (anonymous):

do you perhaps happen to know the formula i should use

hero (hero):

\[a^2 = b^2 + c^2 - \sqrt{2bc \cos A}\]